Grade 11General Physics

A particle is projected vertically upwards from the ground at time t=0 and reaches a height h at t=T. Determine the greatest height reached by the particle

Mayur Reddy
10 Years agoGrade 11

2 Answers

Subhiksha
ApprovedApproved Tutor Answer10 Years ago
Let the particle be projected with a initial velocity u.
 
at time T,it reaches height h.
h=uT-0.5gT2
 
On solving, we get,
u= h/T + gt/2
 
The maximum height reached by the particle id H.
H = u2/2g
Substituting the value of u,
we get,
H= h2/2gT+ gT2/8  + h/2
which is the required equation.
 
Subhiksha
10 Years ago
Just a small error,
“On solving, we get,
u= h/T + gt/2”
It is actually, u= h/T + gT/2
(change in case of T)
Yet, the final answer is right.